Please tell me the exact answer of the que., that how many F-Br-F bond angles are of 90°in BrFPlease tell me the exact answer of the que., that how many F-Br-F bond angles are of 90°in BrFPlease tell me the exact answer of the que., that how many F-Br-F bond angles are of 90°in BrFPlease tell me the exact answer of the que., that how many F-Br-F bond angles are of 90°in BrF5?

Grade:12th pass

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
VSEPR theory says that total no. of electron pairs around Br are 6 as N/2 =(7+5)/2=6
five are bond pairs whereas one is lone pair.
Although structure is octahedral ,there is lone pair-bond pair repulsion between electrons.
So the 4 flourine atoms move towards the upper F atom. So the F-Br-F angle reduces from 90 degrees.
So answer=0.
Hope this helps.
Regards,
Soumya Ranjan Mohanty.